Reports 1-1 of 1 Clear search Modify search
VIS (General)
takafumi.ushiba - 23:18 Friday 13 February 2026 (36363) Print this report
TF measurements for Type-A noise budget

I ran the automeasurement script on the remote desktop session of k1ctr5.
If someone would like to stop the script, please send the command Ctrl+C on the terminal opened on the remote session of k1ctr5.

In this time, I only measured TF that had significant coherence with DARM in the previous measurement (klog35621).

Comments to this report:
takafumi.ushiba - 23:33 Wednesday 18 February 2026 (36391) Print this report

I ran the automeasurement script on the remote desktop session of k1ctr5 to finish the TF measurement for Type-A suspensions.
If someone would like to stop the script, please send the command Ctrl+C on the terminal opened on the remote session of k1ctr5.

Note:

Since excitation flag remains on some models, alignment recoader guardian stops to recoad the IFO alignment.
So I deleted the excitation flag to activate alignment recoader guardian.
Since the current LSC_LOCK guardian needs to update the good oplev values by alignment recoader guardian during lock acquisition, IFO cannot be locked if the excitation flag remains.
So, if  the guardian stops at ALIGNING OMC state for a long time, it would be better to check the excitation flag of all models.

Search Help
×

Warning

×